OpenStudy (dr0zier99):

Match the numerical expressions related to the function f(g) = 4g + 6 with their correct definitions. the output of the function when the input g = 2 -------> the input for the function when the output is 18-----> an input and its corresponding output ---------> the absolute difference of the input and the output when the input is 6-------->

OpenStudy (dr0zier99):

options) 3 24 14 (4,22)

OpenStudy (michele_laino):

hint: if the input is \(g=2\), then the output, is \(f(2)=4 \cdot 2 +6=...?\)

OpenStudy (dr0zier99):

ok i did f(3)=4g+6 and got 18 so 3 is b correct

OpenStudy (dr0zier99):

hope that made sense

OpenStudy (michele_laino):

that's right! if \(g=3\), then \(f(3)=18\)

OpenStudy (dr0zier99):

and f(2)=4⋅2+6 = 14

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (dr0zier99):

so how would i write f(g)=4g+6 with (4,22)???

OpenStudy (michele_laino):

it is simple, since if \(g=4\), then \(f(4)=4 \cdot 4+6=...?\)

OpenStudy (michele_laino):

namely, \(4\) is the input, and \(f(4)=...?\) is the corresponding output

OpenStudy (dr0zier99):

22 (i dont know why but i felt good when i solved that thanks very muvh

OpenStudy (michele_laino):

:)

OpenStudy (dr0zier99):

i wish u were my teacher XD

OpenStudy (michele_laino):

thanks!! :)